# A High-Performance Training-Free Pipeline for Robust Random Telegraph Signal Characterization via Adaptive Wavelet-Based Denoising and Bayesian Digitization Methods

### v1

@@ Methods
In the digitization step, each point in the denoised signal is assigned to discrete amplitude levels detected from KDE peaks, corresponding to distinct multi-trap states (i.e., combinations of individual binary traps). Each continuous-valued amplitude sample is mapped to its nearest KDE peak, effectively converting the signal into a stepwise sequence of discrete levels. These levels represent the most likely trap state configurations at each point in time. The outcomes of this process are illustrated as noisy RTS (gray), denoised RTS (blue), and digitized RTS (red) in Figure 1c-h. The quality of both the denoising and digitization stages can be qualitatively assessed by comparing the digitized trace against the ground truth RTS (black). In the simplest case of a single trap, the digitized signal reduces to a binary time series alternating between two distinct levels, reflecting the “on” and “off” occupancy states. However, as the number of traps increases, the RTS becomes multi-level: a two-trap system, for example, yields four unique amplitude states corresponding to combinations of the two binary trap configurations (00, 01, 10, 11). In such cases, accurate digitization becomes increasingly challenging due to overlapping levels and closely spaced transitions. Several digitization strategies are employed to address this complexity, each with trade-offs in robustness and computational demand. One of the simplest approaches is proximity-based level assignment, extensively used in prior studies [20, 19], where each amplitude point is matched to the closest KDE peak. This method is fast and effective for low-noise, single-trap signals, where transitions are well-separated. However, it becomes error-prone in high-noise or multi-trap contexts, where peak overlap and noise artifacts can lead to misclassification, spurious transitions, and distorted dwell statistics. An HMM-based digitizer was also implemented as a variant during earlier stages of development. The main limitation lies in its reliance on well-initialized transition and emission matrices. Without a strong prior or carefully tuned initialization, the randomly generated starting parameters often lead to unstable convergence, poor accuracy, and high variance in results, especially for multi-trap RTS cases. This makes HMM-based digitization unreliable and unrepresentative of the true signal dynamics, especially when applied in a fully autonomous setting without manual tuning. For these reasons, it is excluded from the formal benchmark comparisons.
@@
Although the DTCWT denoiser is specifically optimized to suppress high-frequency white noise with excellent computational efficiency, it may not fully eliminate residual low-frequency components, such as pink noise, that often persist in RTS signals. To address this, and to further improve upon the capabilities of deep learning-based approaches like the DAEUBL, we introduce a novel Bayesian digitization method designed to complement the DTCWT. In this dual-stage architecture, the DTCWT handles the initial denoising to suppress white noise, while the Bayesian digitizer targets the remaining pink noise, enabling robust state classification across diverse RTS profiles, without neural networks or GPU-intensive processing. The proposed Bayesian digitizer applies probabilistic inference to assign each point in the denoised signal to one of the identified KDE peaks shown in Algorithm 2. Starting with a uniform prior over all peaks, representing equal probability of occupancy, it uses Bayes’ theorem to compute posterior probabilities based on the observed signal value and an estimated noise standard deviation $\sigma$. The peak with the highest posterior is selected as the most likely discrete state. To improve temporal consistency and suppress spurious transitions, the priors are dynamically updated using a smoothing mechanism across time steps. Additionally, an optional refinement step using Dempster-Shafer Theory [31] can be incorporated, leveraging evidence from a short window of previous classifications to further stabilize predictions, especially in multi-trap signals where amplitude levels are closely spaced.
@@
6:  Compute likelihoods: $L_{i}=\exp\left(-\frac{(a(t)-\mu_{i})^{2}}{2\sigma^{2}}\right)/(\sqrt{2\pi}\sigma)$
7:  Compute posteriors: $P_{i}=\frac{\text{prior}_{i}\cdot L_{i}}{\sum_{j}\text{prior}_{j}\cdot L_{j}}$
8:  Assign current value to index $k=\arg\max_{i}P_{i}$
9:  Store $k$ in classification array
10:  Update prior: $\text{prior}\leftarrow\alpha\cdot\text{prior}+(1-\alpha)\cdot P$
11:Initialize output array for binary trap state levels
12:for each classification index $k$ do
13:  Convert index $k$ to binary string of length $N_{\text{trap}}$
14:  Store binary values in output array
15:Save digitized RTS

### v2

@@ Methods
In the digitization step, each point in the denoised signal is assigned to discrete amplitude levels detected from KDE peaks, corresponding to distinct multi-trap states (i.e., combinations of individual binary traps). Each continuous-valued amplitude sample is mapped to its nearest KDE peak, effectively converting the signal into a stepwise sequence of discrete levels. These levels represent the most likely trap state configurations at each point in time. The outcomes of this process are illustrated as noisy RTS (gray), denoised RTS (blue), and digitized RTS (red) in Figure 1c-h. The quality of both the denoising and digitization stages can be qualitatively assessed by comparing the digitized trace against the ground truth RTS (black). In the simplest case of a single trap, the digitized signal reduces to a binary time series alternating between two distinct levels, reflecting the “on” and “off” occupancy states. However, as the number of traps increases, the RTS becomes multi-level: a two-trap system, for example, yields four unique amplitude states corresponding to combinations of the two binary trap configurations (00, 01, 10, 11). In such cases, accurate digitization becomes increasingly challenging due to overlapping levels and closely spaced transitions. Several digitization strategies are employed to address this complexity, each with trade-offs in robustness and computational demand. One of the simplest approaches is proximity-based level assignment, extensively used in prior studies [9, 26], where each amplitude point is matched to the closest KDE peak. This method is fast and effective for low-noise, single-trap signals, where transitions are well-separated. However, it becomes error-prone in high-noise or multi-trap contexts, where peak overlap and noise artifacts can lead to misclassification, spurious transitions, and distorted dwell statistics. An HMM-based digitizer was also implemented as a variant during earlier stages of development. The main limitation lies in its reliance on well-initialized transition and emission matrices. Without a strong prior or carefully tuned initialization, the randomly generated starting parameters often lead to unstable convergence, poor accuracy, and high variance in results, especially for multi-trap RTS cases. This makes HMM-based digitization unreliable and unrepresentative of the true signal dynamics, especially when applied in a fully autonomous setting without manual tuning. For these reasons, it is excluded from the formal benchmark comparisons. For clarity, proximity-based, HMM-based, and Bayesian digitization strategies differ primarily in their sensitivity to residual noise, reliance on parameter initialization, and computational cost: proximity-based methods are fast but noise-sensitive, HMM-based methods are parameter-dependent and unstable without careful initialization, while the proposed Bayesian digitizer offers a lightweight and fully autonomous alternative suited for fluctuating denoised RTS signals.
@@
Although the DTCWT denoiser is specifically optimized to suppress high-frequency white noise with excellent computational efficiency, it may not fully eliminate residual low-frequency components, such as pink noise, that often persist in RTS signals. To address this, and to further improve upon the capabilities of deep learning-based approaches like the DAEUBL, we introduce a novel Bayesian digitization method designed to complement the DTCWT. In this dual-stage architecture, the DTCWT handles the initial denoising to suppress white noise, while the Bayesian digitizer targets the remaining pink noise, enabling robust state classification across diverse RTS profiles, without neural networks or GPU-intensive processing. The proposed Bayesian digitizer applies probabilistic inference to assign each point in the denoised signal to one of the identified KDE peaks shown in Algorithm 2.
@@
6:  Compute likelihoods:
7:   $L_{i}=\exp\left(-\frac{(a(t)-\mu_{i})^{2}}{2\sigma^{2}}\right)/(\sqrt{2\pi}\sigma)$
8:  Compute posteriors: $P_{i}=\frac{\text{prior}_{i}\cdot L_{i}}{\sum_{j}\text{prior}_{j}\cdot L_{j}}$
9:  Assign current value to index $k=\arg\max_{i}P_{i}$
10:  Store $k$ in classification array
11:  Update prior: $\text{prior}\leftarrow\alpha\cdot\text{prior}+(1-\alpha)\cdot P$
12:Initialize output array for binary trap state levels
13:for each classification index $k$ do
14:  Convert index $k$ to binary string of length $N_{\text{trap}}$
15:  Store binary values in output array
16:Save digitized RTS
Bayesian digitization is formulated as a probabilistic inference problem. Let $a(t)\in\mathbb{R}$ denote the denoised RTS amplitude at discrete time index $t=1,\dots,T$. From the denoised signal, we extract a finite set of $K$ discrete amplitude levels $\{\mu_{k}\}_{k=1}^{K}$, obtained as KDE peak locations. We introduce a latent discrete random variable $z_{t}\in\{1,\dots,K\}$, which indicates which KDE level is active at time $t$. We consider the observation model $a(t)=\mu_{z_{t}}+\epsilon_{t}$, where $\epsilon_{t}$ represents the marginal residual noise remaining after denoising. Under this model, digitization corresponds to inferring the most likely discrete state $z_{t}$ given the observed amplitude $a(t)$. We assume that the residual noise $\epsilon_{t}$ is approximately zero-mean Gaussian when considered marginally $p\!\left(a(t)\mid z_{t}=k\right)=\mathcal{N}\!\left(a(t);\mu_{k},\sigma^{2}\right)$. This assumption implies that amplitudes close to $\mu_{k}$ are more likely to arise from state $k$, while deviations are penalized exponentially. The variance $\sigma^{2}$ controls how forgiving the classifier is with respect to deviations from KDE levels. Although RTS noise may exhibit temporal correlation (e.g., pink noise), the Gaussian assumption is applied only to the marginal sample distribution. Temporal stability is handled separately through prior smoothing.
The residual noise standard deviation $\sigma$ is estimated directly from the denoised signal by measuring deviations from the nearest KDE level. For each time step, we define $\hat{\mu}(a(t))=\arg\min_{\mu_{k}}\lvert a(t)-\mu_{k}\rvert$, and compute the root-mean-square residual between the signal and its closest KDE level. Then Bayesian inference is performed at each time step. Let $\pi^{(t)}_{k}\triangleq p(z_{t}=k)$ denote the prior probability of state $k$ at time $t$. At initialization, we use a uniform prior $\pi^{(1)}_{k}=\frac{1}{K}$, reflecting equal occupancy probability across all KDE levels. Given an observation $a(t)$, the likelihood for each state $k$ is computed using a Gaussian emission model $L_{k}=p(a(t)\mid z_{t}=k)$. Applying Bayes’ rule yields the posterior distribution $p(z_{t}=k\mid a(t))$. This update is computed in closed form and requires no iterative optimization. Lastly, a hard digitized label is obtained via maximum a posteriori (MAP) decoding $\hat{z}_{t}=\arg\max_{k}p(z_{t}=k\mid a(t))$. The selected index corresponds to the digitized KDE level at time $t$. To suppress spurious switching caused by residual noise and enforce temporal persistence, we update the prior distribution using exponential smoothing $\pi^{(t+1)}\leftarrow\alpha\,\pi^{(t)}+(1-\alpha)\,p(z_{t}\mid a(t))$, where $0<\alpha<1$. This update acts as a lightweight temporal regularizer; rather than learning a full transition matrix as in HMM, we impose a persistence bias that favors remaining in the same state unless successive observations consistently support a transition to a different dwell time in RTS. Because this update is a convex combination of valid probability distributions, it preserves normalization and non-negativity $\sum_{k}\pi^{(t)}_{k}=1$, $\pi^{(t)}_{k}\geq 0$. Additionally, an optional refinement step using Dempster-Shafer Theory [30] can be incorporated, leveraging evidence from a short window of previous classifications to further stabilize predictions, especially in multi-trap signals where amplitude levels are closely spaced.
The Bayesian digitization method is computationally efficient and scalable. Because the posterior is computed in closed form at each time step, the method avoids iterative optimization as in variational Bayes or Markov chain Monte Carlo. The computational complexity scales linearly $\mathcal{O}(LK)$ with signal length $L$ and number of KDE levels $K$. Compared to HMM-based approaches, which jointly learn emission and transition parameters via expectation-maximization and are sensitive to initialization, our method fixes emission means using data-driven KDE peaks and enforces temporal stability through prior smoothing. Prioritizing robustness, scalability, and fully autonomous analysis of long, high-resolution RTS measurements.
@@
For benchmarking fairness, all methods evaluated in this study are blind to the synthetic benchmark data: none of the denoisers or digitizers are optimized, trained, or tuned on the evaluated RTS signals. The adaptive rules used in the proposed DTCWT denoiser are derived independently from a separate set of real RTS measurements and rely only on general signal statistics, as detailed in Denoising and KDE Algorithm. Baseline methods such as MA and FFT employ adaptive parameters consistent with their standard usage in RTS analysis, while DAEUBL is evaluated using its available pretrained model without retraining. To ensure fair comparison, we adopt a unified and more aggressive noise-scaling scheme across all methods, avoiding the use of the original DAEUBL training data, which would otherwise introduce an unfair advantage.